A car covers its outward journey at 40kmph. It is towed back, over exactly the same distance at 10mph. What is the average speed over the entire journey?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
40 kmph = 40/1.6 mph = 24.85 mph
 
we can apply the formula 2xy/(x+y) for average speed.
Now avg speed = 2*24.85*10/34.85 = 14.26 mph